# Intelligent Design 4.0: Paradigm Evolution Toward the Agentic AI Era

### v1

@@ 2 The Evolution of Intelligent Design Paradigms
Furthermore, as foundation models continue to grow in capacity, the boundary between the multi-agent framework and the single “super-agents” may become blurred [88]. This raises critical questions: At what point could a single model effectively manage the full stack of an end-to-end design workflow? And how should we balance the generality and power of such models against the modularity, interpretability, and fault isolation offered by multi-agent systems? These questions lie at the heart of the orchestration problem in ID 4.0: what is the optimal architecture for intelligent collaboration - distributed agents with explicit communication and specialization, or unified agents with general-purpose capabilities? Future research and practice must explore hybrid architectural strategies that can dynamically evolve with advancements in model capabilities, balancing the strengths of specialization and generality within intelligent design systems.
(4) Toward New Forms of Goal-setting
@@
Engineering design has long been a human-centered activity, with goals explicitly specified by users based on functional requirements, stakeholder needs, and contextual constraints [89]. Even under the ID 4.0 paradigm, human designers typically serve as the ultimate arbiters of what constitutes a desirable outcome [60]. However, as agentic AI systems grow more capable of reasoning, interacting, and learning from context, a new possibility emerges: Can AI agents begin to formulate goals themselves? For example, in response to vague or under-specified prompts, an ID 4.0 system might extrapolate user intent, identify broader system-level opportunities, or suggest higher-level objectives that were previously unconsidered.
The shift from human-defined goals to machine-generated goals may present a fundamental change in how design is conceived and directed [90]. It challenges the long-standing assumption that human intention is the sole source of meaningful goals. If agents are empowered to redefine the design problem or elevate the abstraction level of design targets, they may unlock higher-order creativity and autonomous problem discovery. In doing so, ID 4.0 could evolve from a tool for goal execution into a partner in ideation, capable of augmenting or even co-driving the innovation process.
At the same time, the shift toward machine-defined goals raises fundamental ethical and legal questions [91]. Who is accountable for a machine-generated goal? How do we ensure that such goals respect cultural, environmental, or organizational norms? And what does authorship mean when systems contribute not only as problem-solvers but also as problem-setters?
@@
To better align ID 4.0 systems with human values, we need alignment on two fronts: technically, through transparent reasoning processes and human-in-the-loop validation; and socially, by embedding ethical norms and stakeholder priorities into system design. In addition, relevant policies should ensure that system actions are traceable, responsibilities are well-defined, and human intervention remains possible when needed.
@@ 8 Conclusion
Intelligent Design 4.0 represents a critical inflection point in the ongoing evolution of design intelligence, shifting from data-driven assistance to multi-agent autonomy empowered by foundation AI models. In this paper, we reviewed the paradigm evolution of Intelligent Design through four developmental stages, and presented a multi-agent-based framework for ID 4.0. Beyond reviewing past developments, we have outlined future directions and open challenges toward a more scalable, practical, effective design paradigm. We hope this work offers both conceptual foundations and actionable perspectives to guide researchers, designers, and developers in advancing the next generation of intelligent design systems.

### v2

@@ 7.2 Agent Collaboration Mechanisms
Furthermore, as foundation models continue to grow in capacity, the boundary between the multi-agent framework and the single “super-agents” may become blurred [97]. This raises critical questions: At what point could a single model effectively manage the full stack of an end-to-end design workflow? And how should we balance the generality and power of such models against the modularity, interpretability, and fault isolation offered by multi-agent systems? These questions lie at the heart of the orchestration problem in ID 4.0: what is the optimal architecture for intelligent collaboration, distributed agents with explicit communication and specialization, or unified agents with general-purpose capabilities? Future research and practice should explore hybrid architectural strategies that can dynamically evolve with advancements in model capabilities, balancing the strengths of specialization, negotiation, and generality within the systems.
7.3 Design Problem Formulations
As engineering design problems increase in complexity, they may require integrated reasoning across disparate physical scales (e.g., nano to macro) and physics domains (e.g., structural, thermal, fluid, electromagnetic) [98]. Complex products such as aerospace systems, robotics, and biomedical devices often involve tight coupling between microstructural properties, component-level geometries, control strategies, and performance in dynamic environments [99, 100, 101]. Traditional methods often treat these domains independently, requiring manual iteration across software tools and disciplinary teams. Within the ID 4.0 paradigm, specialized agents could be orchestrated to operate at different modeling scales, enabling coherent and system-level co-design. For instance, in lightweight aerospace component design, a material-level simulation agent could predict anisotropic mechanical properties arising from composite layup patterns, while a structural optimization agent could use these stress-strain and thermal expansion parameters to adjust topology or thickness distributions. The results could then be iteratively evaluated by a manufacturability agent to ensure the optimized structure can be fabricated within process constraints such as curing temperature and fiber orientation tolerance. The development of such cross-domain agents requires close collaboration among experts from different disciplines to ensure accurate and effective results.
Despite this potential, current foundation models still lack accurate embedded knowledge in fields such as nonlinear mechanics, heat transfer, or optics [102]. Their reasoning is largely statistical rather than governed by physical laws. In practice, integrating physics engines or world models with LLMs or VLMs is challenging due to mismatched representations between symbolic solvers and token-based models, the computational cost of fine-scale fidelity, and the need to ensure outputs remain consistent with physical laws. Addressing these gaps will require intermediate representations that bridge symbolic and statistical reasoning [103], as well as coordinated efforts on expanding design data foundations and effective model architectures. Advancing such hybrid systems could ultimately yield cross-scale design agents capable of reasoning from atomic structures to macro-level artifacts [104].
@@ 7.3 Design Problem Formulations
In addition, the formulation of design goals themselves is expected to evolve. Engineering design has traditionally been human-centered, with goals specified by users based on their requirements [105]. However, as agentic AI systems grow more capable of reasoning and contextual learning, they may begin to formulate goals autonomously. For example, when faced with underspecified prompts, a future ID 4.0 system may extrapolate user intent, identify broader opportunities, or suggest higher-level objectives previously unconsidered. This shift challenges the assumption that humans alone define meaningful goals, and positions agents as partners in ideation capable of autonomous problem discovery [106].
@@
At the same time, the shift toward machine-defined goals raises fundamental ethical and legal questions [107]. Who is accountable for a machine-generated goal? How do we ensure that such goals respect cultural, environmental, or organizational norms? And what does authorship mean when systems contribute not only as problem-solvers but also as problem-setters? To better align ID 4.0 systems with human values, we need alignment on two fronts: technically, through transparent reasoning processes and human-in-the-loop validation; and socially, by embedding ethical norms and stakeholder priorities into system design. More importantly, relevant policies should ensure that system actions are traceable, responsibilities are well-defined, and human intervention remains possible when needed.
@@ 8 Conclusion
Intelligent Design 4.0 marks a critical inflection point in the evolution of design intelligence, transitioning from data-driven assistance to multi-agent autonomy empowered by foundation models such as LLMs. In this paper, we reviewed the paradigm evolution of intelligent design through four stages, and proposed a framework for implementing ID 4.0. Beyond reviewing past developments, we have outlined future directions and open challenges, including the need for robust design data foundations, more flexible mechanisms for agent collaboration, and new approaches to formulating multi-scale and evolving design problems and goals. We hope this paper could offer both conceptual foundations and actionable perspectives to guide researchers, designers, and developers in advancing the next generation of intelligent design.